If I know P(A^B), how do I get P(A^Bcomplement)...where ^ means intersection ....for example
P(A) = P(B)_= 1/3 and P(A^B) = to 1/10
What is (a) P(A U Bcomplement)
(b) P(B^Acomplement)
(c) P(Acomplement U Bcomplement)

Answer by stanbon(75887)

If I know P(A^B), how do I get P(A^Bcomplement)...where ^ means intersection ....
--------------
P(A) + P(A comp) = 1
So, P(A comp) = 1-P(A)
--------------------------
P(A^B comp) = 1 - P(A^B)
--------------------------
for example
P(A) = P(B)_= 1/3 and P(A^B) = to 1/10
What is (a) P(A U Bcomplement)
Find P(AUB)
P(AUB) = P(A)+P(B)-P(A^B)
= 1/3 + 1/3 - 1/10
= (2/3)-(1/10)
= (20-3)/30 = 17/30
---
The P(AUB comp) = 13/30
=============================
(b) P(B^Acomplement) = P(B-(A^B))
= P(B)-P(A^B)
= P(B)-[P(A)+P(B)-P(AUB)]
= P(AUB)-P(A)
= (17/30)-(1/3)
= 7/30
===============
(c) P(Acomplement U Bcomplement)
= 1-P(A^B)
= 1-[P(A)+P(B)-P(AUB)]
= 1-[1/3 + 1/3 - (17/30)]
= 1-[(20/30)-(17/30)]
= 1 - 1/10
= 9/10
================
Cheers,
Stan H.
================